# PhysLab: A Benchmark Dataset for Multi-Granularity Visual Parsing of Physics Experiments

### v1

@@ 4. Selected Tasks
With its diverse and multi-granularity annotations, the PhysLab dataset supports a wide array of visual analysis tasks, including temporal action localization, action classification, action recognition, object detection, occlusion detection, instance segmentation, HOI detection, and scene graph generation. In this section, we evaluate the performance of several representative methods on PhysLab, focusing on two key tasks: action recognition and HOI detection. These tasks are used to examine the unique challenges of visual parsing in physics experiment scenarios from both temporal and spatial perspectives. For additional tasks and benchmarks, please refer to the official open-source repository.
@@ 4.1. Action Recognition
Task Formulation. The objective of action recognition is to locate and classify actions from untrimed videos. In the context of procedural task analysis, this task is typically divided into two subtasks (Li et al., 2019; Lu and Elhamifar, 2021):
@@
(1) $\frac{1}{T}\sum_{i=1}^{T}(\hat{y_{i}}=y_{i})$
(2) $\frac{1}{A}\sum_{a}\left|G_{a}\cap D_{a}\right|/\left|G_{a}\cup D_{a}\right|$
@@
PhysLab HICO-DET
Method Backbone Full Rare Non-Rare Full Rare Non-Rare
STIP (Zhang et al., 2022b) ResNet-50 62.62 61.00 62.73 32.22 28.15 33.43
GEN-VLKTs (Liao et al., 2022) ResNet-50 58.71 75.00 57.58 33.75 29.25 35.10
OCN (Yuan et al., 2022) ResNet-50 52.19 68.01 51.10 30.91 25.26 32.51
LOGICHOI (Li et al., 2023) ResNet-50 53.34 50.97 62.50 35.47 32.03 36.22
TED-Net (Wang et al., 2024a) ResNet-50 60.97 69.89 60.35 34.00 29.88 35.24
SOV-STG-S (Chen et al., 2025) ResNet-50 45.42 34.17 49.50 33.80 29.28 35.15
OCN (Yuan et al., 2022) ResNet-101 49.50 50.00 49.46 31.43 25.80 33.11
UPT (Zhang et al., 2022a) ResNet-101 65.28 63.47 65.43 32.31 28.55 33.44
GEN-VLKTl (Liao et al., 2022) ResNet-101 60.19 64.25 69.91 34.95 31.18 36.08
SOV-STG-L (Chen et al., 2025) ResNet-101 52.35 46.59 58.44 35.01 30.63 36.32
ERNet (Lim et al., 2023) ENetV2-XL 56.39 49.26 62.36 35.92 30.13 38.29
PViC (Zhang et al., 2023) Swin-L 71.42 68.47 71.62 44.32 44.61 44.24
FGAHOI (Ma et al., 2024) Swin-L 60.79 54.89 67.23 37.18 20.71 39.11
SOV-STG-Swin-L (Chen et al., 2025) Swin-L 56.48 45.23 58.44 43.35 42.25 43.69
Results. We evaluated four representative methods across three datasets—PhysLab, Breakfast, and CrossTask—covering both action alignment and action segmentation tasks. As summarized in Table 2, these methods show overall inferior performance on PhysLab compared to the other two datasets, particularly in terms of MoF. This performance gap can be attributed to several factors inherent to PhysLab: the presence of numerous fine-grained experimental actions, frequent occlusions, complex human-object interactions, and high visual similarity between distinct operations. These characteristics substantially increase the difficulty of both frame-level action classification and temporal boundary localization.
Moreover, PhysLab exhibits a stronger capacity for distinguishing model performance. For the action alignment task, the IoU gap between the best and worst-performing models on PhysLab reaches 21.2, whereas the corresponding gaps on Breakfast and CrossTask are only 5.0 and 3.1, respectively. A similar trend is observed in the action segmentation task, where PhysLab presents an IoU gap of 20.3, again notably larger than those on the other datasets. These results highlight that the complex visual and procedural characteristics of PhysLab not only introduce greater challenges for action recognition but also provide a more rigorous benchmark for assessing model robustness and generalization capabilities.
@@ 4.2. Human-Object Interaction Detection
Evaluation Metrics. We adopt the standard mean Average Precision (mAP) under the Full, Rare, and Non-Rare settings, following the HICO-DET evaluation protocols (Chao et al., 2018). A predicted interaction triplet is considered true positive (TP) if and only if all of the following conditions are met:
@@
Results. We evaluated 10 representative HOI detection methods on PhysLab and HICO-DET, with the results summarized in Table 3. Overall, model performance on PhysLab is generally higher than on HICO-DET, primarily because PhysLab focuses on structured and domain-specific interactions in physics experiments, whereas HICO-DET covers diverse and open-ended scenarios. Despite this performance advantage, PhysLab introduces two unique challenges. First, its pronounced category imbalance results in significant performance shifts across classes. Unlike HICO-DET, where most models consistently perform better on Non-rare categories, models on PhysLab exhibit divergent trends. For instance, STIP (Zhang et al., 2022b) and LOGICHOI (Li et al., 2023) perform better on Non-rare interactions, while OCN (Yuan et al., 2022) achieves superior performance on Rare categories (achieving 68.01 mAP on Rare classes vs. 51.10 on Non-rare). Notably, GEN-VLKTs (Liao et al., 2022) exhibits a 17.42 mAP gap between two sets, highlighting the increased generalization demands posed by PhysLab under imbalanced label distributions.
Second, model performance exhibits greater volatility across models in each evaluation subset. Specifically, within the Full, Rare, and Non-Rare subsets of PhysLab, the performance gap between the best and worst models reaches 57.2%, 119.5%, and 44.8% respectively, which are significantly higher than the corresponding gaps observed in HICO-DET, at 43.4%, 115.4%, and 20.3%. This suggests that PhysLab’s interaction scenarios not only amplify the strengths and weaknesses of different model architectures, but also better reveal their varying adaptability to distinct interaction types and category frequencies. Therefore, the pronounced class imbalance and greater performance disparities in PhysLab make it a more challenging benchmark (Chan et al., 2024). These characteristics underscore its value in promoting the development of more robust and fine-grained HOI detection algorithms, particularly in experimental or instructional contexts (Hu et al., 2024).
@@ 5. Conclusion and Future work
In this paper, we introduce PhysLab, a richly annotated benchmark dataset designed to advance multi-granularity visual parsing of physical experiment processes. PhysLab fills a critical void at the intersection of computer vision and science education by embedding visual recognition tasks within authentic experimental contexts. By modeling the complex procedural structures inherent in experimental activities, PhysLab facilitates the development of interpretable and goal-oriented scene understanding systems tailored to educational environments. We believe that PhysLab offers a novel and impactful perspective for both educational AI research and the broader vision community.
We are actively expanding the dataset by collecting and annotating video data for six additional representative physics experiments. In parallel, we plan to extend PhysLab to encompass experimental procedures in chemistry and biology, thereby increasing domain diversity and supporting cross-disciplinary generalization. In addition, to enable more robust and context-aware modeling of experimental processes, we aim to integrate multimodal signals such as audio, textual instructions, and sensor data, and explore cross-modal alignment and fusion strategies to better capture the semantic structure of experimental tasks. All versions of the dataset, along with benchmark protocols and evaluation results, are continuously maintained and released through our open-source platform to encourage reproducibility and foster collaborative development within the research community.

### v2

@@ 4. Selected Tasks
As summarized in Table 1, PhysLab offers rich, diverse, and multi-granularity annotations, enabling a wide range of visual analysis tasks. In this section, we benchmark two representative settings with complementary emphases: video-level action recognition, which evaluates a model’s capacity to leverage temporal context, and image-level HOI detection, which assesses spatial reasoning in complex experimental environments. This dual-focus evaluation is intended to highlight the distinct challenges of visual parsing in physics experiment scenarios from both temporal and spatial perspectives. For additional tasks and benchmarks, please refer to the official open-source repository.
@@ 4.1. Action Recognition
Task Formulation. The objective of action recognition is to locate and classify actions from untrimmed videos. In the context of procedural analysis (Li et al., 2019; Lu and Elhamifar, 2021), this task is divided into two subtasks:
@@
(1) $\frac{1}{T}\sum_{i=1}^{T}(\hat{y_{i}}=y_{i}),$
(2) $\frac{1}{A}\sum_{a}\left|G_{a}\cap D_{a}\right|/\left|G_{a}\cup D_{a}\right|,$
@@
Results. We evaluated four representative methods across three datasets: PhysLab, Breakfast, and CrossTask, with the latter two selected for being the authoritative and widely adopted procedural video benchmarks. As summarized in Table 2, these methods show overall inferior performance on PhysLab compared to the other two datasets, particularly in terms of MoF. This performance gap can be attributed to several factors inherent to PhysLab: the presence of diverse experimental actions, frequent occlusions, complex human-object interactions, and high visual similarity between distinct operations. These characteristics substantially increase the difficulty of both frame-level action classification and temporal boundary localization.
Moreover, PhysLab exhibits a stronger capacity for distinguishing model performance. For the action alignment task, the IoU gap between the best and worst-performing models on PhysLab reaches 21.2, whereas the corresponding gaps on Breakfast and CrossTask are only 5.0 and 3.1, respectively. A similar trend is observed in the action segmentation task, where PhysLab presents an IoU gap of 20.3, again notably larger than those on the other datasets. These results highlight that the complex visual and procedural characteristics of PhysLab not only introduce greater challenges for action recognition, but also provide a more rigorous benchmark for assessing model robustness and generalization capabilities.
@@ 4.2. Human-Object Interaction Detection
Evaluation Metrics. We adopt the standard mean Average Precision (mAP) under the Full, Rare, and Non-Rare settings, following the HICO-DET evaluation protocols (Chao et al., 2018). A predicted interaction triplet is considered a true positive (TP) if and only if all of the following conditions are met:
@@
Results. We evaluated 10 representative HOI detection methods on PhysLab and HICO-DET, adopting the latter as the reference benchmark since other datasets listed in Table 1 lack HOI annotations. Results are summarized in Table 3. Following (Liao et al., 2022; Chen et al., 2025), the backbones are organized into three groups: ResNet-50, ResNet-101, and Swin-L. Overall, model performance on PhysLab is generally higher than on HICO-DET, primarily because PhysLab focuses on structured and domain-specific interactions in physics experiments, whereas HICO-DET covers diverse and open-ended scenarios. Despite this performance advantage, PhysLab introduces two unique challenges. First, its pronounced category imbalance results in significant performance shifts across classes. Unlike HICO-DET, where most models consistently perform better on Non-rare categories, models on PhysLab exhibit divergent trends. For instance, STIP (Zhang et al., 2022b) and LOGICHOI (Li et al., 2023) perform better on Non-rare interactions, while OCN (Yuan et al., 2022) achieves superior performance on Rare categories (achieving 68.01 mAP on Rare classes vs. 51.10 on Non-rare). Notably, GEN-VLKTs (Liao et al., 2022) exhibits a 17.42 mAP gap between the two sets, highlighting the increased generalization demands posed by PhysLab under imbalanced label distributions.
PhysLab HICO-DET
Method Full Rare Non-Rare Full Rare Non-Rare
STIP (Zhang et al., 2022b) 62.62 61.00 62.73 32.22 28.15 33.43
GEN-VLKTs (Liao et al., 2022) 58.71 75.00 57.58 33.75 29.25 35.10
OCN (Yuan et al., 2022) 52.19 68.01 51.10 30.91 25.26 32.51
LOGICHOI (Li et al., 2023) 53.34 50.97 62.50 35.47 32.03 36.22
TED-Net (Wang et al., 2024a) 60.97 69.89 60.35 34.00 29.88 35.24
SOV-STG-S (Chen et al., 2025) 45.42 34.17 49.50 33.80 29.28 35.15
OCN (Yuan et al., 2022) 49.50 50.00 49.46 31.43 25.80 33.11
UPT (Zhang et al., 2022a) 65.28 63.47 65.43 32.31 28.55 33.44
GEN-VLKTl (Liao et al., 2022) 60.19 64.25 69.91 34.95 31.18 36.08
SOV-STG-L (Chen et al., 2025) 52.35 46.59 58.44 35.01 30.63 36.32
PViC (Zhang et al., 2023) 71.42 68.47 71.62 44.32 44.61 44.24
FGAHOI (Ma et al., 2024) 60.79 54.89 67.23 37.18 20.71 39.11
SOV-STG-Swin-L (Chen et al., 2025) 56.48 45.23 58.44 43.35 42.25 43.69
Second, model performance exhibits greater relative volatility across models in each evaluation subset. Specifically, within the Full, Rare, and Non-Rare subsets of PhysLab, the performance gap between the best and worst models (calculated as the percentage difference normalized by the worst model’s score) reaches 57.2%, 119.5%, and 44.8%, respectively. These gaps are significantly higher than those observed in HICO-DET, which are 43.4%, 115.4%, and 20.3%, respectively. This suggests that PhysLab’s interaction scenarios not only amplify the strengths and weaknesses of different model architectures, but also better reveal their varying adaptability to distinct interaction types and category frequencies (Chao et al., 2018). As described in (Liu et al., 2019; Chan et al., 2024), the greater inter-class performance differences in PhysLab make it a more challenging benchmark (Chan et al., 2024). These characteristics underscore PhysLab’s value in promoting the development of more robust and advanced HOI detection algorithms (Zou et al., 2021), particularly in experimental or instructional contexts (Hu et al., 2024).
@@ 5. Conclusion and Future Work
In this paper, we introduce PhysLab, a richly annotated dataset designed to advance multi-granularity visual parsing of physical experiment processes. PhysLab fills a critical void at the intersection of computer vision, multimedia, and science education by embedding visual recognition tasks within authentic experimental contexts.
We will actively expand the dataset by annotating data for six additional representative physics experiments. Concurrently, we plan to collect experimental procedures in chemistry and biology, thereby increasing domain diversity, enabling cross-domain adaptation tests, and supporting cross-disciplinary generalization. To enable more robust and context-aware modeling of experimental processes, we aim to integrate multimodal signals such as audio, textual instructions, and sensor data, and explore cross-modal alignment and fusion strategies to better capture the semantic structure of experimental tasks. All versions of the dataset, along with benchmark protocols and evaluation results, will be continuously maintained and released through our open-source platform.
Acknowledgements.
This work was supported in part by the National Major Science and Technology Projects of China award [2022ZD0119501], National Natural Science Foundation of China award [52374221], and Science and Technology Development Fund of Shandong Province of China award [ZR2023MF097].